# Mohammad Mohaiminul Islam 1   Rishabh Anand $^{2\textsuperscript{*}}$   David R. Wessels $^{4\textsuperscript{*}}$

### v1

@@ 3.1 Weight-sharing across RoPE Embeddings
Scores for each frame are computed in parallel as their own independent attention head. Note that we can also obtain $s_{ij}(R)$ from the perspective of steering the base set of frequencies $\Omega$ instead of the positions $\mathbf{p}_{i}$, which we show in Appendix E. However, from our current perspective, the RoPE-attention mechanism itself remains completely unchanged from its traditional formulation in Eq. 4; only the relative positions $\mathbf{p}_{i}-\mathbf{p}_{j}$ are now defined relative to each reference frame $R$. The attention coefficients are obtained by applying the softmax to the scores $s_{ij}(R)$. The output $\mathbf{y}_{i}(R)$ for each input feature $\mathbf{f}_{i}(R)$ is then given as
@@ 3.2 Equivariant Linear Layers and Fixed Computation Graph
Crucially, by choosing the number of channels $C$ such that the effective feature dimension $C\times|\mathcal{G}|$ is held constant, the overall matrix dimensions are identical regardless of the group size. The trivial group $\mathcal{G}=\{e\}$ illustrates the base case, where the operation collapses to a standard linear layer with a weight matrix $\mathbf{W}_{\text{group}}(e)$ of size $C^{\prime}\times C$. The geometric inductive bias is therefore not introduced by adding new, complex modules, but by imposing a structure on the weights of existing ones. [^1] This structure can even give computational benefits, by implementing the linear layers in the Fourier domain of $\mathcal{G}$ (Bökman et al., 2025). In Appendix M, we find that at marginally higher channel counts than used in this paper, a Fourier implementation leads to greatly improved training throughput, indicating that this is a promising direction for future research.
With all components of the architecture now defined as equivariant operations, we can formally state the key property of the full model, proven in detail in Appendix C.
@@ 4.1 Platonic Transformer as Dynamic Group Convolution
The use of RoPE in a linear attention setting establishes a deep connection to convolution. Specifically, the mechanism implements an adaptive convolution where the kernel is synthesized on-the-fly. This dynamic kernel is expressed as an expansion in a sparse Fourier basis, defined by the RoPE frequencies, and the coefficients for this basis expansion are provided by the query vectors. This makes the convolution content-aware. We formalize this as follows (proof in Appendix D.1).
@@
Within our Platonic Transformer, this entire mechanism is lifted to operate over the reference frames defined by a group $\mathcal{G}$. Consequently, the operator becomes an adaptive group convolution (proof in Appendix D.3), where the kernel is steered by the group elements/reference frames.
@@ 5.1 Experimental Setup
For fair comparison, we match the computational cost between $SE(n)$ and $T(n)$ models by equating our group-based parallelism with standard multi-head attention. For instance, an $SE(n)$ model using the 12-element tetrahedral group with one head per frame is benchmarked against a $T(n)$ baseline with 12 total heads (details in App. H). For certain tasks, symmetries can be conditionally broken by using APE or providing an external reference frame, yet internal layers critically retain principled weight-sharing. Using external frames to break symmetry and APE to provide geometric information are effective strategies, allowing a model to benefit from geometric inputs without being end-to-end constrained by full equivariance (Vadgama et al., 2025).
For CIFAR-10 and ScanObjectNN, we conducted a comprehensive sweep to find the optimal configuration. In contrast, for QM9 and OMol25, we used a sequential process: first, we identified the best architecture via an extensive sweep on QM9, then transferred these hyperparameters to OMol25 for further refinement with a one-million subset before full training (see Appendix I-L).
@@ 5.3 Equivariant Tasks
Our discussion on the use of reference frames from platonic solid symmetry groups is also reminiscent of Frame Averaging (FA) (Puny et al., 2022), a framework that incorporates group equivariance via symmetrization (detailed introduction in Appendix F) to non-equivariant neural network backbones. FA can be extended to full $E(3)$ by including roto-reflections (e.g., roto-reflecting PCA axes) but comes at a cost, requiring a separate forward pass for each frame element. Our Platonic Transformer offers a more scalable drop-in replacement. For instance, using just a single orthogonal matrix from PCA (i.e., just one frame element) along with an axis-aligned roto-reflection group (here, we use $C_{2}\times C_{2}\times C_{2}$, a subgroup of the octahedral group) achieves native $E(3)$ equivariance at an $\sim 8$x lower cost. We evaluate the effectiveness of FA by comparing to FAFormer (Huang et al., 2024), which applies FA to a standard Transformer backbone with 8 frame elements. Our single-frame variant with only octahedral reflections (8-Refl + 1 frame) outperforms FAFormer with a smaller compute cost. This underscores the need for a geometrically expressive backbone, such as the Platonic Transformer.
@@ 6 Conclusion
We make several strides towards reproducibility of our work. We back our theoretical results with proofs in Appendix D, and with rough intuitions in the main text. We have attached a .zip folder as supplementary material with source code to train our Platonic Transformer models (and baselines) across different tasks and datasets; we also provide adequate comments and documentation. Details required to reproduce the results in our tables and figures are provided in Appendices I-L. We also provide information on hyperparameter-tuning efforts in Appendices G and H. We intend to open-source the code on public platforms like GitHub once the review period has formally ended.
@@ A.4 A Fourier Perspective
Appendix B Related Works
Equivariant Transformers.
SE(3) Transformer, EFA, G-Hyena, EScAIP, GATr
Approximate Equivariance.
Frame Averaging, REMUL, EquiGrad in Orb-v3, SymPE
Appendix C Equivariance Properties of Platonic Transformers
@@
Appendix D Proofs
D.1 Proof of Proposition 2
@@ D.2 Proof of Corollary 1
D.3 Proof of Platonic Transformers Implementing Group Convolutions
@@
Appendix E Equivalent Attention via RoPE Base Frequency Steering
@@
Appendix F Frame Averaging
@@
Appendix G Details of Architecture
@@
Appendix H Hyperparameter Tuning and Model Selection Strategy
@@
Appendix I Details of experiments on Cifar10
I.1 Description of the dataset
@@ I.2 Training Details
Appendix J Details of experiments on ScanObjectNN
J.1 Description of the dataset
@@
Appendix K Details of experiments on QM9
K.1 Description of the dataset
@@
K.2 Training Details for the Regression Experiment
@@
K.3 Additional details on wall-clock timings
@@
Appendix L Details of experiments on OMol25
L.1 Description of the dataset
@@
Appendix M Implementing Platonic Transformers in the Fourier Domain of Finite Groups
@@
M.1 Introduction to the Fourier Theory of Finite Groups
@@
The features in Platonic Transformers are functions from $\mathcal{G}$ to $\mathbb{R}^{C}$, that transform under the left regular representation as explained in Appendix C. In order words, the representation that acts on them is a direct sum of $C$ copies of the regular representation of $\mathcal{G}$. Let this representation be denoted $\tilde{\rho}$. Decomposing $\tilde{\rho}$ into irreps, we obtain
@@
M.2 Fourier Theory of the Tetrahedral Group
@@
Appendix N Disclosure of LLM Usage

### v2

@@ 3.1 Weight-sharing across RoPE Embeddings
Scores for each frame are computed in parallel as their own independent attention head. Note that we can also obtain $s_{ij}(R)$ from the perspective of steering the base set of frequencies $\Omega$ instead of the positions $\mathbf{p}_{i}$, which we show in Appendix D. However, from our current perspective, the RoPE-attention mechanism itself remains completely unchanged from its traditional formulation in Eq. 4; only the relative positions $\mathbf{p}_{i}-\mathbf{p}_{j}$ are now defined relative to each reference frame $R$. The attention coefficients are obtained by applying the softmax to the scores $s_{ij}(R)$. The output $\mathbf{y}_{i}(R)$ for each input feature $\mathbf{f}_{i}(R)$ is then given as
@@ 3.2 Equivariant Linear Layers and Fixed Computation Graph
Crucially, by choosing the number of channels $C$ such that the effective feature dimension $C\times|\mathcal{G}|$ is held constant, the overall matrix dimensions are identical regardless of the group size. The trivial group $\mathcal{G}=\{e\}$ illustrates the base case, where the operation collapses to a standard linear layer with a weight matrix $\mathbf{W}_{\text{group}}(e)$ of size $C^{\prime}\times C$. The geometric inductive bias is therefore not introduced by adding new, complex modules, but by imposing a structure on the weights of existing ones. [^1] This structure can even give computational benefits, by implementing the linear layers in the Fourier domain of $\mathcal{G}$ (Bökman et al., 2025). In Appendix L, we find that at marginally higher channel counts than used in this paper, a Fourier implementation leads to greatly improved training throughput, indicating that this is a promising direction for future research.
With all components of the architecture now defined as equivariant operations, we can formally state the key property of the full model, proven in detail in Appendix B.
@@ 4.1 Platonic Transformer as Dynamic Group Convolution
The use of RoPE in a linear attention setting establishes a deep connection to convolution. Specifically, the mechanism implements an adaptive convolution where the kernel is synthesized on-the-fly. This dynamic kernel is expressed as an expansion in a sparse Fourier basis, defined by the RoPE frequencies, and the coefficients for this basis expansion are provided by the query vectors. This makes the convolution content-aware. We formalize this as follows (proof in Appendix C.1).
@@
Within our Platonic Transformer, this entire mechanism is lifted to operate over the reference frames defined by a group $\mathcal{G}$. Consequently, the operator becomes an adaptive group convolution (proof in Appendix C.3), where the kernel is steered by the group elements/reference frames.
@@ 5.1 Experimental Setup
For fair comparison, we match the computational cost between $SE(n)$ and $T(n)$ models by equating our group-based parallelism with standard multi-head attention. For instance, an $SE(n)$ model using the 12-element tetrahedral group with one head per frame is benchmarked against a $T(n)$ baseline with 12 total heads (details in App. G). For certain tasks, symmetries can be conditionally broken by using APE or providing an external reference frame, yet internal layers critically retain principled weight-sharing. Using external frames to break symmetry and APE to provide geometric information are effective strategies, allowing a model to benefit from geometric inputs without being end-to-end constrained by full equivariance (Vadgama et al., 2025).
For CIFAR-10 and ScanObjectNN, we conducted a comprehensive sweep to find the optimal configuration. In contrast, for QM9 and OMol25, we used a sequential process: first, we identified the best architecture via an extensive sweep on QM9, then transferred these hyperparameters to OMol25 for further refinement with a one-million subset before full training (see Appendix H-K).
@@ 5.3 Equivariant Tasks
Our discussion on the use of reference frames from platonic solid symmetry groups is also reminiscent of Frame Averaging (FA) (Puny et al., 2022), a framework that incorporates group equivariance via symmetrization (detailed introduction in Appendix E) to non-equivariant neural network backbones. FA can be extended to full $E(3)$ by including roto-reflections (e.g., roto-reflecting PCA axes) but comes at a cost, requiring a separate forward pass for each frame element. Our Platonic Transformer offers a more scalable drop-in replacement. For instance, using just a single orthogonal matrix from PCA (i.e., just one frame element) along with an axis-aligned roto-reflection group (here, we use $C_{2}\times C_{2}\times C_{2}$, a subgroup of the octahedral group) achieves native $E(3)$ equivariance at an $\sim 8$x lower cost. We evaluate the effectiveness of FA by comparing to FAFormer (Huang et al., 2024), which applies FA to a standard Transformer backbone with 8 frame elements. Our single-frame variant with only octahedral reflections (8-Refl + 1 frame) outperforms FAFormer with a smaller compute cost. This underscores the need for a geometrically expressive backbone, such as the Platonic Transformer.
@@ 6 Conclusion
We make several strides towards reproducibility of our work. We back our theoretical results with proofs in Appendix C, and with rough intuitions in the main text. We have attached a .zip folder as supplementary material with source code to train our Platonic Transformer models (and baselines) across different tasks and datasets; we also provide adequate comments and documentation. Details required to reproduce the results in our tables and figures are provided in Appendices H-K. We also provide information on hyperparameter-tuning efforts in Appendices F and G. We intend to open-source the code on public platforms like GitHub once the review period has formally ended.
@@ A.4 A Fourier Perspective
Appendix B Equivariance Properties of Platonic Transformers
@@
Appendix C Proofs
C.1 Proof of Proposition 2
@@ C.2 Proof of Corollary 1
C.3 Proof of Platonic Transformers Implementing Group Convolutions
@@
Appendix D Equivalent Attention via RoPE Base Frequency Steering
@@
Appendix E Frame Averaging
@@
Appendix F Details of Architecture
@@
Appendix G Hyperparameter Tuning and Model Selection Strategy
@@
Appendix H Details of experiments on Cifar10
H.1 Description of the dataset
@@ H.2 Training Details
Appendix I Details of experiments on ScanObjectNN
I.1 Description of the dataset
@@ I.2 Training Details
Appendix J Details of experiments on QM9
J.1 Description of the dataset
@@
J.2 Training Details for the Regression Experiment
@@
J.3 Additional details on wall-clock timings
@@
Appendix K Details of experiments on OMol25
K.1 Description of the dataset
@@
Appendix L Implementing Platonic Transformers in the Fourier Domain of Finite Groups
@@
L.1 Introduction to the Fourier Theory of Finite Groups
@@
The features in Platonic Transformers are functions from $\mathcal{G}$ to $\mathbb{R}^{C}$, that transform under the left regular representation as explained in Appendix B. In order words, the representation that acts on them is a direct sum of $C$ copies of the regular representation of $\mathcal{G}$. Let this representation be denoted $\tilde{\rho}$. Decomposing $\tilde{\rho}$ into irreps, we obtain
@@
L.2 Fourier Theory of the Tetrahedral Group
@@
Appendix M Disclosure of LLM Usage